OpenStudy (anonymous):

in a triangle prove that bcos^2C/2+ccos^2B/2=s here s is semiperimeter of a triangle

OpenStudy (anonymous):

Is it (bcos2c)/2

OpenStudy (anonymous):

U know half angle theorems right Use those equations to find the solution I got it by that method

OpenStudy (anonymous):

As cos C/2=square root of (s(s-c))/ab. Cos B/2=square root of (s(s-b))/ac

OpenStudy (anonymous):

From numerator and denominator b and c get cancelled

OpenStudy (anonymous):

Then take s/a com on to get s/a(s-b+s-c) =s/a(2s-b-c) = (s/a)*a = s hence proved

OpenStudy (anonymous):

thank u i did some mistake in the last step

OpenStudy (anonymous):

Ur welcome

OpenStudy (anonymous):

Any more problems